| <p>11:30</p>
<p>Run 50 started.</p>
<p>D+ beam deuterium target. D+ beam energy is 0.35 MeV/u. (beam calc calculation is 0.3577 MeV/u)</p>
<p>Target density is about 4e12 at/cm2</p>
<p>All four detectors positioned fully out and stationary.</p>
<p>Left potentiometer = 18.86 mm</p>
<p>Right potentiometer = 19.8 mm</p>
<p>Ring cycle time is 10 seconds.</p>
<p>Target T1=30.1K, T2=31.8K.</p>
<p>Nozzle position unchanged since end of run 48</p>
<p>YR09-Etarget pressure = 2.05e-10 mbar.</p>
<p>Cooler Read value = 207.1 V, 4.39 mA</p>
<p> </p>
<p>All leakage currents look good (except dssd4, currently stable at V=60)</p>
<p>Beam position. Vertical =0.0mm , Horizontal=0.5 mm</p>
<p>Data rate ~ 172kB/s</p>
<p>Slow comparator threshold to 0x14 (200 keV) since we decreased beam energy.</p>
<p>13:00</p>
<p>All checks OK</p>
<p>We were seeing some rubbish obscuring the 3He peak. JM remembered setting the EYEX cut very high (3 MeV) on the online plot, Now it is reduced to 500 keV</p>
<p>See attachments of histograms before restarting code.<br />
</p>
<p>16:32 all checks okay<br />
18:19 all checks okay<br />
20:20 all checks okay<br />
21:45 all checks okay</p>
<p>01:00 all checks okay</p>
<p>04:00 all checks okay</p>
<p> </p>
<p>7:00</p>
<p>Run stopped - cave to be opened. Beam looks on when run was stopped.</p>
<p>Temps OK - attach</p>
<p>Leakage current stable - attach</p>
<p>Rates OK - attach</p>
<p>Merger OK - attach</p>
<p>Final spectra attached</p>
<p>09:30</p>
<p>Gas target shutter valve closed to save ring vacuum.</p>
<p>All checks OK</p>
<p>Stats at htis threshold with no beam attached.</p>
<p>Biased DSSD4 in steps of 5V from 60V since it has been stable since yesterday,</p>
<p>Was able to bias to 100V with no issues.</p> |